OpenStudy (anonymous):

an alligator with a body length of 4.5 feet has a tail of 3.5 feet. What is the tail length of an alligator whose body length is 3.6 feet

OpenStudy (anonymous):

4.5/3.5 = 3.6/x 4.5x=12.6 x=2.8 feet

OpenStudy (anonymous):

3.6*3.5=4.5t 12.6 = 4.5t t= 12.6/4.5 t=2.8 ft